# Overestimation, Overfitting, and Plasticity in Actor-Critic: the Bitter Lesson of Reinforcement Learning

### v1

@@ Abstract
Recent advancements in off-policy Reinforcement Learning (RL) have significantly improved sample efficiency, primarily due to the incorporation of various forms of regularization that enable more gradient update steps than traditional agents. However, many of these techniques have been tested in limited settings, often on tasks from single simulation benchmarks and against well-known algorithms rather than a range of regularization approaches. This limits our understanding of the specific mechanisms driving RL improvements. To address this, we implemented over 60 different off-policy agents, each integrating established regularization techniques from recent state-of-the-art algorithms. We tested these agents across 14 diverse tasks from 2 simulation benchmarks. Our findings reveal that while the effectiveness of a specific regularization setup varies with the task, certain combinations consistently demonstrate robust and superior performance. Notably, a simple Soft Actor-Critic agent, appropriately regularized, reliably solves dog tasks, which were previously solved mainly through model-based approaches.
@@ 1 Introduction
Despite significant advancements, the understanding of how different regularization techniques synergistically improve off-policy agent performance is still limited (Hiraoka et al., 2021; Lee et al., 2023). Moreover, most methods are tested in narrow contexts, mainly in locomotion or manipulation tasks, often restricted to a single simulation benchmark (Fujimoto et al., 2018; Haarnoja et al., 2018; Chen et al., 2020; Moskovitz et al., 2021; D’Oro et al., 2022), leading to questions about their broad applicability and robustness. In this study, our goal is to consolidate these lessons and address the following research questions: Which regularization techniques lead to robust performance improvements across diverse tasks and agent designs? Can generic regularization techniques outperform domain-specific RL techniques that directly use the MDP structure?. We extend the scope of prior research by examining over 60 design choices in a diverse array of tasks, including both locomotion and manipulation, within two simulation benchmarks and two replay ratio regimes. This comprehensive approach offers a deeper understanding of the effectiveness of these regularization techniques in various settings.
Our main result reinforces the ’Bitter Lesson’ (Sutton, 2019): across varied tasks, general neural network regularizers significantly outperform most RL-specific algorithmic improvements in terms of agent performance. Notably, network regularization enables agents to solve tasks previously impossible for model-free agents, such as those in the dog domain. Our findings also show that layer normalisation is more effective in reducing overestimation than techniques specifically designed for mitigating Q-value overestimation in critic networks. Consequently, we show that replacing the ubiquitous Clipped Double Q-learning with network regularization techniques leads to significant performance gains. Our research further explores the impact of overestimation, overfitting, and plasticity loss on agent performance in a unified experimental setup. We examine the correlation between these factors and agent performance, showing a strong negative correlation for value overestimation and agent plasticity metrics. These influences vary significantly across environments, underscoring the complex nature of their effects on learning. A key observation is the environment-dependent performance of various methods. Strategies excelling in locomotion tasks may falter in manipulation scenarios and vice versa. Comparisons between experiments on the DeepMind Control Suite (Tassa et al., 2018) and MetaWorld (Yu et al., 2019) demonstrate the necessity for diverse benchmarking in research, highlighting the value of expansive experimental setups.
@@ 2.1 Overestimation
Q-learning methods employing function approximation have been observed to exhibit a bias toward overestimation, a phenomenon critical to the training process (Thrun & Schwartz, 2014; Fujimoto et al., 2018). Positive bias stems from the policy being trained to locally maximize action-value estimates, leading its actions to exploit potential model errors for higher scores. Modern actor-critic algorithms leverage a variety of countermeasures to overestimation of Q-value targets, with Clipped Double Q-learning (CDQ) (Fujimoto et al., 2018) being most used by many other algorithms (Haarnoja et al., 2018; Chen et al., 2020; Hiraoka et al., 2021). In CDQ, the algorithm maintains an two critics and uses their minimum as an approximate a Q-value lower bound. The CDQ was generalized to the following pessimistic objective (Ciosek et al., 2019; Moskovitz et al., 2021; Cetin & Celiktutan, 2023):
@@ 2.2 Overfitting
Overfitting, while not commonly scrutinized in reinforcement learning, has gained attention in recent discussions (Li et al., 2022) as a phenomenon correlated with performance decline in models characterized by a high ratio of updates to data. To evaluate overfitting in agents, Li et al. (2022) utilize a validation dataset that consists of samples gathered using the same policy as the canonical replay buffer. The validation buffer is established to provide an unbiased assessment of the critic error on experiences that were not used in the learning. Although there are many strategies to deal with overfitting in supervised learning, only few of them were applied in the context of RL. To this end, application of Weight Decay (WD) (Schwarzer et al., 2023), Layer Normalization (LN) (Ball et al., 2023) or Spectral Normalization (SN) (Cetin & Celiktutan, 2023) was shown to greatly effect the performance of the underlying agent.
@@ 3 Study Design
    • Layer Norm (Ba et al., 2016),
    • Spectral Norm (Miyato et al., 2018; Zhang et al., 2018; Brock et al., 2018),
    • Weight Decay (Loshchilov & Hutter, 2017),
@@
Third-order results involve no intervention marginalization and represent the performance of a specific combination, including one method from each group. The only marginalization is over all tested environments (we present these results in Appendix D.1 due to space constraints). This cumulative result provides insight into the overall impact of a given intervention.
Furthermore, we conduct an analysis of various proxy metrics associated with the problems of overestimation, overfitting and plasticity loss. For overestimation, we evaluate the state-action critic approximation error, denoted as $b_{\phi}(s,a)$, is quantified as the disparity between the critic output and the true on-policy Q-value according to $b_{\phi}(s,a)=Q_{\phi}(s,a)-Q^{\pi}(s,a)$, where $Q_{\phi}$ denotes the critic Q-value approximation and $Q^{\pi}$ represents the on-policy Q-value which we estimate via a Monte-Carlo rollout with 5 samples. To calculate overfitting, we compare average TD errors on evaluation trajectories (which are not used for learning) to average TD errors observed in training according to $o_{\phi}=\frac{\mathbb{E}_{\mathcal{D}_{v}}TD_{\phi}}{\mathbb{E}_{\mathcal{D}}TD_{\phi}}$, where $o_{\phi}$ denotes critic overfitting, $\mathcal{D}_{v}$ denotes validation data, and $TD_{\phi}$ denotes the temporal difference loss. As such, the extent of overestimation is then quantified by the ratio of validation TD error to training TD error. We monitor plasticity loss by the rank of penultimate layer representations (Kumar et al., 2020), dormant neurons or dead units (Sokar et al., 2023), and the L2 norm of weights as a proxy for plasticity loss.
@@ 4.1 Combination of interventions – First-order marginalization
First-order marginalization provides insights into the robust impact of a given intervention on model performance, irrespective of what other type if regularization it is paired with. To measure such robustness, we compare the performance of the baseline SAC model augmented with one specific regularization (e.g., SAC + WD) to the performance of SAC augmented this regularization paired with some other technique (e.g. SAC + WD + Resets).
@@ Results:
Examining the plots in Figure 1 shows that network and plasticity regularization techniques are generally more effective than critic regularization. We observe this results for both simple models using one regularization technique and more complex agents leveraging many regularizations at once. Most notably, avoiding the use of critic regularization interventions (!CR) proves advantageous for both DMC and MW, especially if some other type of regularization is used (such as layer norm or full-parameter resets). This result is somewhat surprising, as critic regularization methods where designed specifically for off-policy actor-critic agents, whereas the network and plasticity regularization techniques are general. Notably, TOP (Moskovitz et al., 2021) emerges as an exception, particularly showcasing its effectiveness on the DMC benchmark. Conversely, the GPL intervention exhibits the least robust performance in both DMC and MW. Full-parameter resets (Nikishin et al., 2022), tailored for high replay ratio regimes, proves to be one of the most robust approaches in this RR regime. Further analysis reveals discrepancies in conclusions between benchmarks. Clearly, LN is the most effective approach on the DMC benchmark, yet it presents very poor performance on the MW benchmark. To this end, we find SN to be more robust, which strongly helps the MW benchmark and moderately helps the DMC.
@@
Critic regularization methods exhibit limited effectiveness in enhancing performance. When using network or plasticity regularization, critic regularization leads to reduced performance. Periodical network resetting is the most robust intervention across two benchmarks in a high RR regime, and highly surpasses other plasticity regularization techniques in both robustness and performance. When considering network regularization approaches, layer norm is generally recommended for DMC, while layer norm is more effective for MW benchmarks. When considering a diverse tasks, we find spectral norm to be more robust than layer norm. Weight decay has generally low performance when used alone with SAC.
Figure 1: IQM performance of First-Order Marginalization. The left column presents results for baseline SAC augmented with a single regularization technique (and thus uses 10 seeds per task), and the right column presents the aggregate performance of a specific regularization technique when paired with other regularizations (and thus uses 640 seeds per task). Results are presented for DMC (top row) and MW (bottom row) benchmarks. 14 tasks.
@@ 4.2 Combination of interventions – Second-order marginalization
Figure 2: Second-order results marginalizing critic regularization methods. On the x-axis, we have different types of plasticity regularization, and each color denotes network regularization. For better readability, points within one plasticity regularization are spaced slightly horizontally. Vertical lines indicate standard error.
@@ Results:
On the DMC with RR=2 and RR=16 (top row in Figure 2) a clear hierarchy of interventions is observed: layer norm and right below it, spectral norm consistently outperforms others in mean return, irrespective of the plasticity regularization (x-axis). Notably, the combination of layer norm and resets in RR=16 (top-right plot in Figure 2) demonstrates exceptional performance across all critic regularization variations. In contrast, the hierarchy of interventions on the MW benchmark (bottom row in Figure 2) diverges significantly from the DMC setup. Furthermore, a higher replay ratio introduces shifts in training dynamics, as evidenced by SN transitioning from a lower position on RR=2 (bottom-left plot in Figure 2) to nearly the most versatile intervention on RR=16 (bottom-right plot in Figure 2). The results generally align with first-order marginalization findings, emphasizing the positive impact of SN, as well as using many different types of regularization at once in general. Deeper analysis (see Appendix D.2) reveals that on MW, indeed, the gradient norm in a higher RR regime is orders of magnitude bigger. The finding that most contrasts with the first-order experiments, is that we observe that weight decay can actually yield significant performance benefits, under the condition that it is paired with other specific method, namely full-parameter resets. In particular, we observe that this combination yields synergies surpassing using any of these methods alone.
@@
4.3 A closer look on Dog environment
@@
In this study, we delve into the intricacies of two challenging Dog tasks, Dog-Trot and Dog-Run, included in our DMC setup. These environments present considerable difficulties for model-free approaches relying on proprioceptive states, making them of particular interest within the research community. For detailed experimental information, please refer to Appendix B.
@@ Results:
Figure 3: IQM performance evolution across timesteps for Dog-Run (top row) and Dog-Trot (bottom row) environments. Each plot showcases the top four combinations alongside vanilla SAC (coded as: CDQ+!NR+!PR).
Specific intervention combinations effectively tackle the challenges posed by the Dog environment, as depicted in Figure 3. Analyzing the top four approaches for Dog-Run and Dog-Trot tasks reveals the prevalence of layer norm in nearly all combinations. Additionally, each critic regularization approach contributes to the leading group. Notably, in scenarios with high replay ratios, resets emerge as a crucial intervention. These observations are further substantiated by the analysis of second-order marginalization IQM plots (see 4). Indeed, layer norm without critic regularization excels in RR=2, and layer norm with resets outperforms all others convincingly.
This achievement is particularly notable as, to our best knowledge, no model-free agent has previously solved the Dog environment using proprioceptive states.
Figure 4: IQM performance of the top six intervention pair combinations. The IQM is calculated based on the average of the last ten evaluation points in each run, not the last evaluation point. Top row: Dog-Run environment. Bottom row: Dog-Trot. Colors indicate hierarchy on the plot, not specific names.
@@
Well-established network regularization techniques such as layer norm and spectral norm enable solving Dog-Trot and Dog-Run effectively. The choice of domain-specific RL critic regularization has little significance in dog environments when layer norm and resetting interventions are employed.

### v2

@@ Abstract
Recent advancements in off-policy Reinforcement Learning (RL) have significantly improved sample efficiency, primarily due to the incorporation of various forms of regularization that enable more gradient update steps than traditional agents. However, many of these techniques have been tested in limited settings, often on tasks from single simulation benchmarks and against well-known algorithms rather than a range of regularization approaches. This limits our understanding of the specific mechanisms driving RL improvements. To address this, we implemented over 60 different off-policy agents, each integrating established regularization techniques from recent state-of-the-art algorithms. We tested these agents across 14 diverse tasks from 2 simulation benchmarks, measuring training metrics related to overestimation, overfitting, and plasticity loss — issues that motivate the examined regularization techniques. Our findings reveal that while the effectiveness of a specific regularization setup varies with the task, certain combinations consistently demonstrate robust and superior performance. Notably, a simple Soft Actor-Critic agent, appropriately regularized, reliably finds a better-performing policy within the training regime, which previously was achieved mainly through model-based approaches.
@@ 1 Introduction
Despite significant advancements, the understanding of how different regularization techniques synergistically improve off-policy agent performance is still limited (Hiraoka et al., 2021; Lee et al., 2023). Moreover, most methods are tested in narrow contexts, mainly in locomotion or manipulation tasks, often restricted to a single simulation benchmark (Fujimoto et al., 2018; Haarnoja et al., 2018; Chen et al., 2020; Moskovitz et al., 2021; D’Oro et al., 2022), leading to questions about their broad applicability and robustness. In this study, our goal is to consolidate these lessons and address the following research questions: Which regularization techniques lead to robust performance improvements across diverse tasks and agent designs? Can generic regularization techniques outperform domain-specific RL techniques that directly use the MDP structure?. We extend the scope of prior research by examining over 60 design choices implemented within the Soft Actor-Critic framework. We test a diverse array of tasks, including both locomotion and manipulation, within two simulation benchmarks and two replay ratio regimes. This comprehensive approach offers a deeper understanding of the effectiveness of these regularization techniques in various settings.
Our main result is a bitter lesson: across varied tasks, general neural network regularizers significantly outperform most RL-specific algorithmic improvements in terms of agent performance. Specifically, we find general methods that are motivated by stabilization of gradient-based learning significantly outperform RL-specific algorithmic improvements across a variety of environments. Such emphasis on generality is in line with the celebrated “Bitter Lesson” essay (Sutton, 2019). Notably, network regularization enables agents to find effective policies on tasks previously impossible for model-free agents, such as those in the dog domain. Our findings also show that layer normalisation is more effective in reducing overestimation than techniques specifically designed for mitigating Q-value overestimation in critic networks. Consequently, we show that replacing the ubiquitous Clipped Double Q-learning with network regularization techniques leads to significant performance gains. Our research further explores the impact of overestimation, overfitting, and plasticity loss on agent performance in a unified experimental setup. We examine the correlation between these factors and agent performance, showing a strong negative correlation for value overestimation and agent plasticity metrics. These influences vary significantly across environments, underscoring the complex nature of their effects on learning. A key observation is the environment-dependent performance of various methods. Strategies excelling in locomotion tasks may falter in manipulation scenarios and vice versa. Comparisons between experiments on the DeepMind Control Suite (Tassa et al., 2018) and MetaWorld (Yu et al., 2019) demonstrate the necessity for diverse benchmarking in research, highlighting the value of expansive experimental setups.
@@ 2.1 Overestimation
Q-learning methods employing function approximation have been observed to exhibit a bias toward overestimation, a phenomenon critical to the training process (Thrun & Schwartz, 2014; Fujimoto et al., 2018). Positive bias stems from the policy being trained to locally maximize action-value estimates, leading its actions to exploit potential model errors for higher scores. Modern actor-critic algorithms leverage a variety of countermeasures to overestimation of Q-value targets, with Clipped Double Q-learning (CDQ) (Fujimoto et al., 2018) being most used by many other algorithms (Haarnoja et al., 2018; Chen et al., 2020; Hiraoka et al., 2021). In CDQ, the algorithm maintains two critics and uses their minimum as an approximate Q-value lower bound. The CDQ was generalized to the following pessimistic objective (Ciosek et al., 2019; Moskovitz et al., 2021; Cetin & Celiktutan, 2023):
@@ 2.2 Overfitting
Overfitting, while not commonly scrutinized in reinforcement learning, has gained attention in recent discussions (Li et al., 2022) as a phenomenon correlated with performance decline in models characterized by a high ratio of updates to data. To evaluate overfitting in agents, Li et al. (2022) utilizes a validation dataset that consists of samples gathered using the same policy as the canonical replay buffer. The validation buffer is established to provide an unbiased assessment of the critic error in experiences that were not used in the learning. Although there are many strategies to deal with overfitting in supervised learning, only a few of them were applied in the context of RL. To this end, application of Weight Decay (WD) (Schwarzer et al., 2023), Layer Normalization (LN) (Ball et al., 2023) or Spectral Normalization (SN) (Cetin & Celiktutan, 2023) was shown to greatly effect the performance of the underlying agent.
@@ 3 Study Design
    • Layer Norm (LN) (Ba et al., 2016),
    • Spectral Norm (SN) (Miyato et al., 2018; Zhang et al., 2018; Brock et al., 2018),
    • Weight Decay (WD) (Loshchilov & Hutter, 2017),
@@
Third-order results involve no intervention marginalization and represent the performance of a specific combination, including one method from each group. The only marginalization is over all tested environments (we present these results in Appendix C.1 due to space constraints). This cumulative result provides insight into the overall impact of a given intervention.
Furthermore, we conduct an analysis of various proxy metrics associated with the problems of overestimation, overfitting, and plasticity loss. For overestimation, we evaluate the state-action critic approximation error, denoted as $b_{\phi}(s,a)$, is quantified as the disparity between the critic output and the true on-policy Q-value according to $b_{\phi}(s,a)=Q_{\phi}(s,a)-Q^{\pi}(s,a)$, where $Q_{\phi}$ denotes the critic Q-value approximation and $Q^{\pi}$ represents the on-policy Q-value which we estimate via a Monte-Carlo rollout with 5 samples. To calculate overfitting, we compare average TD errors on evaluation trajectories (which are not used for learning) to average TD errors observed in training according to $o_{\phi}=\frac{\mathbb{E}_{\mathcal{D}_{v}}TD_{\phi}}{\mathbb{E}_{\mathcal{D}}TD_{\phi}}$, where $o_{\phi}$ denotes critic overfitting, $\mathcal{D}_{v}$ denotes validation data, and $TD_{\phi}$ denotes the temporal difference loss. As such, the extent of overestimation is then quantified by the ratio of validation TD error to training TD error. We monitor plasticity loss by the rank of penultimate layer representations (Kumar et al., 2020), dormant neurons or dead units (Sokar et al., 2023), the L2 norm of weights (Nikishin et al., 2022; Lyle et al., 2023), and gradient norm (Nikishin et al., 2022; Lyle et al., 2023) as a proxy for plasticity loss.
@@ 4.1 Combination of interventions – First-order marginalization
First-order marginalization provides insights into the robust impact of a given intervention on model performance, irrespective of what other type of regularization it is paired with. To measure such robustness, we compare the performance of the baseline SAC model augmented with one specific regularization (e.g., SAC + WD) to the performance of SAC augmented with this regularization paired with some other technique (e.g. SAC + WD + Resets).
Figure 1: IQM performance of First-Order Marginalization. The left column presents results for baseline SAC augmented with a single regularization technique (and thus uses 10 seeds per task), and the right column presents the aggregate performance of a specific regularization technique when paired with other regularizations (and thus uses 640 seeds per task). Results are presented for MW (top row), DMC without Dog environments (middle row) and only Dog-run and Dog-trot (bottom row) benchmarks. 14 tasks.
@@ Results:
Examining the plots in Figure 1 shows that network and plasticity regularization techniques are generally more effective than critic regularization – dark blue line (!CR) on Robustness plots on MW and DMC (!DOC). We observe these results for both simple models using one regularization technique and more complex agents leveraging many regularizations at once. Most notably, avoiding the use of critic regularization interventions (!CR) proves advantageous for both DMC and MW, especially if some other type of regularization is used (such as layer norm or full-parameter resets). This result is somewhat surprising, as critic regularization methods were designed specifically for off-policy actor-critic agents, whereas the network and plasticity regularization techniques are general. Notably, TOP (Moskovitz et al., 2021) emerges as an exception, particularly showcasing its effectiveness on the DMC benchmark. Conversely, the GPL intervention exhibits the least robust performance in both DMC and MW. Upon further analysis of the impact of CDQL presence (see Section C.4), it becomes apparent that in certain environments, such as Hopper Hop, the adverse effects of this intervention cannot be mitigated even with additional regularizations. Full-parameter resets (Nikishin et al., 2022), tailored for high replay ratio regimes, prove to be one of the most robust approaches in this RR regime. Further analysis reveals discrepancies in conclusions between benchmarks. Clearly, LN is the most effective approach in the DMC Dog environments (bottom row in Figure 1), but it also ranks among the top four interventions in the remaining DMC environments. However, it exhibits very poor performance on the MW benchmark. Therefore, we find SN to be more robust, significantly aiding the MW benchmark and providing moderate assistance in the DMC scenarios.
@@
Critic regularization methods exhibit limited effectiveness in enhancing performance. When using network or plasticity regularization, critic regularization leads to reduced performance. Periodical network resetting is the most robust intervention across two benchmarks in a high replay ratio regime, and highly surpasses other plasticity regularization techniques in both robustness and performance. Layer norm is essential for Dog environments. When considering network regularization approaches, layer norm is generally recommended for DMC, while spectral norm is more effective for MW benchmarks. When considering a diverse range of tasks, we find spectral norm to be more robust than layer norm. Weight decay has generally low performance when used alone with SAC.
@@ 4.2 Combination of interventions – Second-order marginalization
Figure 2: Second-order results marginalizing critic regularization methods. On the x-axis, we have different types of plasticity regularization, and each colour denotes network regularization. For better readability, points within one plasticity regularization are spaced slightly horizontally. Vertical lines indicate standard error.
@@ Results:
On the DMC with RR=2 and RR=16 (top row in Figure 2) a clear hierarchy of interventions is observed: layer norm and right below it, spectral norm consistently outperforms others in mean return, irrespective of the plasticity regularization (x-axis). Notably, the combination of layer norm and resets in RR=16 (middle and bottom-right plot in Figure 2) demonstrates exceptional performance across all critic regularization variations on DMC Dog and without Dog environments. In contrast, the hierarchy of interventions on the MW benchmark (top row in Figure 2) diverges significantly from the DMC setup. Furthermore, a higher replay ratio introduces shifts in training dynamics, as evidenced by SN transitioning from a lower position on RR=2 (bottom-left plot in Figure 2) to nearly the most versatile intervention on RR=16 (bottom-right plot in Figure 2). The results generally align with first-order marginalization findings, emphasizing the positive impact of SN, as well as using many different types of regularization at once in general. Deeper analysis (see Appendix C.2) reveals that on MW, indeed, the gradient norm in a higher RR regime is orders of magnitude bigger. The finding that most contrasts with the first-order experiments, is that we observe that weight decay can actually yield significant performance benefits, under the condition that it is paired with other specific methods, namely full-parameter resets. In particular, we observe that this combination yields synergies surpassing using any of these methods alone.
@@
4.3 A closer look on Dog environment performance
@@
In this study, we delve into the intricacies of two challenging Dog tasks, Dog-Trot and Dog-Run, included in our DMC setup. These environments present considerable difficulties for model-free approaches relying on proprioceptive states, making them of particular interest within the research community. Due to the inherent difficulty of these tasks, we conducted additional experiments using the top three methods identified in Figure 23 (Appendix) for 4 million steps, akin to approaches used in model-based (Hansen et al., 2022) or pixel-based studies (Ji et al., 2024). For the rest of the detailed experimental information, please refer to Appendix A.
@@ Results:
Figure 3: Mean return evolution across 4 million timesteps for Dog-Run (top row) and Dog-Trot (bottom row) environments. Gray plot depicts model-based agent performance. Each plot showcases the top three combinations.
Specific intervention combinations effectively tackle the challenges posed by the Dog environment, as depicted in Figure 3. Analyzing the top three approaches for Dog-Run and Dog-Trot tasks reveals the prevalence of layer norm in nearly all combinations. Additionally, each critic regularization approach contributes to the leading group. Notably, in scenarios with high replay ratios, resets emerge as a crucial intervention. These observations are further substantiated by the analysis of second-order marginalization IQM plots (see 4). Indeed, layer norm without critic regularization excels in RR=2, and layer norm with resets outperforms all others convincingly. This achievement is particularly notable as, to our best knowledge, no model-free agent has previously find a better-performing policy within the training regime the Dog environments using proprioceptive states. Notably, there is a recent study(Ji et al., 2024) where a model-free agent achieved comparable results on the Dog environments but using pixel-based inputs instead. Additionally, our results demonstrate that while a model-based approach on proprioceptive states (Hansen et al., 2022) outperforms slightly, the above model-free approach with simple regularization techniques achieves performance very close to that of the model-based approach. This suggests the efficacy and competitiveness of our approach in challenging environments.
Figure 4: IQM performance of the top six intervention pair combinations based on 1 million steps experiments. The IQM is calculated based on the average of the last ten evaluation points in each run, not the last evaluation point. Results come from 1 million steps experiments. Top row: Dog-Run. Bottom row: Dog-Trot.
@@
Well-established network regularization techniques such as layer norm and spectral norm enable finding high-performing policies for Dog-Trot and Dog-Run effectively. The choice of domain-specific RL critic regularization has little significance in dog environments when layer norm and resetting interventions are employed.